OpenStudy (grant330sims):

Find the Direction Angle: u=-4i-7j

OpenStudy (anonymous):

To find the Direction Angle, we only have to do: \[\theta=arctan\left(\frac{y}{x}\right)\] \[\theta=arctan\left(\frac{-7}{-4}\right)\] \[\theta\approx60.25^o\approx1.05~rad\]

OpenStudy (grant330sims):

You back at it again DEX! lol